substitution F2,-)) 13,6] y = 2x - 4 6x - 3y = 12 y = 2x - 4 6x - 3y = 12 son, 2012
Answer #1

We can solve this by substitution method.9+2= 0) --- siblitute 1 in 2 246 +3230 2 120+ 6+30 3x+9=0 3 3x+3=0 x -3 subtitute to So the past (x, y) = (-3, 3) 5x-y = -5 -22- 34 - 12-0 x+y=9 - y=9-2 subtitute in 6 X-3 (9-x) = -12 2x-27 + 3x = -12 52 - 27 = - 5x = -12 +27 substitute in 3+y = 9 Yx=-3 - 0 X-34 = -15 - substitute in a -3-3y = 15 --34 = - 15+3_ -34 = -12 4 = -1213 (6,y) = (-3, 4)4 = 22- 4 0 Gx-3y = 12 -2 solve an equation for a subtitute in 2x = 4+4 x= yty 6 (974) – 3y & 12 by + 24 - 3y = 12 12 Y + d 3substitute in y = 2x - 4 -4 =2x-4 22 - 4+4 2x=0 074